OpenStudy (anonymous):

need help with algebra word problem one number is 3 times as large as another. The sum of their reciprocals is 20/3. find the two numbers.

OpenStudy (mathstudent55):

Can you pick a variable for the smaller number?

OpenStudy (anonymous):

i honestly have no clue on how to do this >_<

OpenStudy (mathstudent55):

Let's call the smaller number x. The other number is 3 times larger, so it's 3x. Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Do you know what a reciprocal is?

OpenStudy (anonymous):

is it when you flip the numbers?

OpenStudy (mathstudent55):

Correct. Now we have two numbers, x and 3x. When you add their reciprocals you get 20/3. The reciprocal of x is \(\dfrac{1}{x}\). The recipoprocal of 3x is \( \dfrac{1}{3x} \). Now we add the reciprocals and set that equal to \( \dfrac{20}{3} \).

OpenStudy (mathstudent55):

That means we have \( \dfrac{1}{x} + \dfrac{1}{3x} = \dfrac{20}{3} \) Now we need to solve for x.

OpenStudy (anonymous):

how would i go about doing that? I'm sorry, math is really not my best subject

OpenStudy (mathstudent55):

Since you have an equation with fractions, the way to get rid of all denominators is to multiply both sides by the LCD. What is the LCD of x, 3x, and 3?

OpenStudy (anonymous):

would it be 3x? or 3?

OpenStudy (mathstudent55):

3x is correct

OpenStudy (mathstudent55):

Now we multiply both sides by 3x \(3x \left( \dfrac{1}{x} + \dfrac{1}{3x} \right) = 3x \cdot \dfrac{20}{3} \) \( \ 3x \cdot \dfrac{1}{x} + 3x \cdot \dfrac{1}{3x} = 3x \cdot \dfrac{20}{3} \) \( \ 3\cancel x \cdot \dfrac{1}{\cancel x} + \cancel{3x} \cdot \dfrac{1}{\cancel{3x}} = \cancel 3x \cdot \dfrac{20}{\cancel{3}} \) \(3 + 1 = 20x \) Can you solve for x now?

OpenStudy (anonymous):

x=5?

OpenStudy (mathstudent55):

Be careful. \(4 = 20x\) Divide both sides by 20 \( \dfrac{4}{20} = x\) \( x = \dfrac{1}{5} \) The smaller number is 1/5. The larger number is 3 times larger, so it's 3/5. The numbers are 1/5 and 3/5.

OpenStudy (mathstudent55):

Now let's check. 3/5 is 3 times large than 1/5, so the larger number is indeed 3 times larger than the smaller number. Now let's add the reciprocals of 1/5 and 3/5 to see if we get 20/3. 5 + 5/3 = 15/3 + 5/3 = 20/3 The sum checks out, so the answer is correct.

OpenStudy (anonymous):

i always have the hardest time with word problems, but thank you so much for your help!

OpenStudy (mathstudent55):

You're welcome. To translate a word problem into an equation, select a variable and follow the statements.